# AADNet: An End-to-End Deep Learning Model for Auditory Attention Decoding

### v1

@@ 3.2.2 Audio
The audio envelopes were extracted using the compressed subband envelopes, resembling the processing of speech signals in the human auditory system [11]. Specifically, we first applied a gammatone filter bank with an equivalent rectangular bandwidth (ERB) equal to 1.5 Hz, and center frequencies ranging from 150 Hz to 4 kHz. The compressed envelope of each subband was computed with a power law exponent of 0.6. The final envelope was then obtained by summing all the subband envelopes. For the NSR model, the envelopes were extracted by taking the absolute value of the Hilbert transform of each speech stream. The envelopes were then low-pass filtered up to 50 Hz using FIR filter, Hamming window, 12.5 Hz transition bandwidth, and resampled to 125 Hz. The gammatone filter bank was designed using the COCOHA Matlab toolbox [33]) while both envelope extraction methods were implemented in Python using the SciPy package.
@@ 3.3 Evaluation procedure
As pointed out in Rotaru et al.[34], the DL approach is susceptible to capturing subtle biases, such as within-trial fingerprints of neural activities even across different subjects. This may lead to artificially high decoding accuracies. To mitigate this potential bias, it is essential to perform appropriate data splitting into training, validation, and test sets. In this study, we carefully employ a trial-based cross validation to evaluate the SS and SI models for all investigated methods. In the remainder of this paper, we refer to the length of data used by the models to make a decision as the analysis window length.
@@ 3.3.2 SI models
For the SI model, data from one subject were held out for the test set while the data from the remaining subjects were used for training and validation. However, the EventAAD dataset was collected using the same pairs and order of stimuli for all subjects which could lead to potential leakage of stimuli information from the training set to the test set, allowing the model to indirectly learn to perform the AAD task. This is not the case for the DTU dataset where the order and pair of stimuli were randomly mixed. To ensure that the testing data were not from any trial that contained training and/or validation data, we performed an eight-fold leave-one-subject-out (LOSO) cross-validation scheme. The data from the test subject were divided into eight folds. Only one fold was held out for the test set while the other seven folds were not used in the current iteration. For all the remaining training subjects, the single held-out test fold was left out of the training set, and trials that were not in the test fold were pooled together across subjects and split into a training set and a validation set in a 4:1 ratio (based on a trial basis). Models were trained using the training and validation set and evaluated on the test set to obtain a fold performance. In this manner, the trial used for the test subject was not used during training, even for the other subjects. See Figure 4 for a visual illustration of the procedure. This procedure was repeated across eight folds. The accuracy of each SI model was calculated as the average accuracy across the eight folds.
@@ 3.4.2 Minimal expected switch duration
The accuracy metric described above is a window length-dependent quantification. It is expected that the longer the window length, i.e., the more information available, the higher the accuracy. To obtain a more effective performance measurement, we also calculated the minimal expected switch duration (MESD), an interpretable performance metric for AAD algorithms in a context of neuro-steered gain control [35]. The MESD addresses the trade-off between AAD accuracy and decision time by modeling an adaptive gain control system in a hearing device as a Markov chain, and based on that calculating the minimal expected time required to switch the operation mode after an attention switch of the user. A lower MESD corresponds to better AAD performance and vice versa. In this study, the MESD was calculated using the Python MESD toolbox [36].
@@ 3.5 Hyperparameter choice and model training
The hyper-parameters of the linear methods were selected based on the recommended values used in the original studies as discussed in Section 2. To determine the optimal $J$ value in the CCA method, we first performed inner cross-validation on the training data for the LDA model to find the optimal $J_{f}$ for each fold from the outer cross-validation of the CCA model. The final $J$ values for SS and SI models were obtained by taking the average and grand average of the $J_{f}$, respectively. These final $J$ values were then used to validate the corresponding CCA models again on the test data to obtain the final performances. For the NSR model, we used the recommended hyper-parameters and settings from the original study and tuned the learning rate.
The proposed AADNet was trained with the Cross-Entropy loss function using the AdamW optimizer [37] which is advantageous in decoupling L2 regularization, i.e., weight decay, from the learning rate so that we have less number of hyper-parameters to search. The hyperparameters were chosen via a random search within configurations as follows: batch size = (32, 64, 128), weight decay = ($10^{-4}$, $10^{-3}$, $10^{-2}$, $10^{-1}$), dropout = (0.5, 0.4, 0.3, 0.2, 0.1) and number of output channel of FC1 layer, i.e., hidden units, $h$ = (32, 16, 0). In the case of $h$ = 0, the ELU, DO, BN, and FC2 after the FC1 layer were omitted. The learning rate was fixed at $10^{-5}$. During the training process, the model was saved if the validation loss decreased and the training process was stopped if the validation loss did not decrease in at least 1 out of 5 consecutive epochs. Although AADNet is flexible regarding input window length, we trained different models per window length to maximize the performance. Due to limited data sets for SS models that potentially lead to overfitting, for the CCA, NSR and the proposed AADNet models, we started training the SI models first and used the saved model to fine-tune the SS models for the subject that was left out. We only fine-tuned the very last FC layers (LDA model for the CCA method) with smaller tuned learning rates.
@@ 4.1 SS models
Decoding accuracies of the SS models for each method on the EventAAD and DTU datasets are shown in Figure 5 and Figure 5, respectively. The box plots represent the accuracy distribution across 24 subjects for the EventAAD dataset and 18 subjects for the DTU dataset. The chance performance was computed as the $95^{th}$ percentile point of a binomial distribution with $p=0.5$ and $n$ equal to the number of non-overlapping windows in the test set. We compared the difference between the proposed AADNet and the other baseline methods and tested the significance using the paired permutation test [40] with the Bonferroni correction. The tested results are shown at the bottom of the figures. On the EventAAD dataset, the mean accuracy increases from 0.573 at 1 s up to 0.802 at 20 s. For the DTU dataset, mean accuracies of 0.619 and 0.88 were obtained at window lengths of 1 s and 40 s, respectively. Notably, it is observed that the AADNet model significantly outperforms all baseline methods across all window lengths on the EventAAD dataset ($p<0.01$), and short window lengths (up to 5 s) on the DTU dataset.
Figure 5: Comparison of the SS models on the two datasets. (a), (b) The accuracies and MESD of the four methods on the EventAAD dataset. (c), (d) The accuracies and MESD of the four methods on the DTU dataset. The number of data points with an MESD of $>100$ s is indicated as (+x) and were included in the computation of the medians. Statistical significance is denoted by asterisks (None: $p\geq 0.05$; *: $0.05>p\geq 0.01$; **: $0.01>p\geq 0.001$; ***: $0.001>p$.
Additionally, we calculated the per subject MESD values for each model on both datasets and showed them in Figure 5 and Figure 5, respectively. The significant differences in the median values between the proposed model and the other were also tested using paired permutation tests. We found that the proposed AADNet generates significantly lower MESD values (23.0 s on the EventAAD dataset and 11.8 s on the DTU dataset) compared to other models.
@@ 4.2 SI models
To evaluate how well AADNet generalizes to new subjects, we performed the SI validation scheme described in Section 3.3.2. The results of SI models are shown in Figure 6. Similar to the SS results, Figure 6 and Figure 6 present the test accuracies across subjects while Figure 6 and Figure 6 present the MESDs. Generally, the SI models obtained a consistently lower performance compared to the corresponding SS models (see Table 3). The results demonstrate that AADNet significantly outperforms the other methods for all window lengths on the EventAAD dataset with mean accuracies reaching up to 0.785 at 20 s. For the DTU dataset, EEGNet obtains significantly higher accuracies for window lengths greater than 2 s and reaches 0.827 at 40 s. At 1 s, the AADNet and CCA models obtained a similar performance with accuracies of 0.561, and 0.562, respectively. Regarding MESD, AADNet achieves significantly lower values with 31.7 s and 29.2 s on the EventAAD and DTU datasets, respectively.
Figure 6: Comparison of the SI models on the two datasets. (a), (b) The accuracies and MESD of the four methods on the EventAAD dataset. (c), (d) The accuracies and MESD of the four methods on the DTU dataset.
Table 3: Performance drop (in percentage points) of SI models compared to corresponding SS models.
@@
EventAAD LSR 0.1 0.5 0.6 1.0 1.6
CCA 0.5 1.1 1.4 1.7 0.8
NSR 0.2 0.2 0.2 0.6 0.5
AADNet 0.5 1.2 1.9 1.1 1.9
DTU LSR 4.0 6.0 8.8 10.7 10.5 11.7
CCA 2.0 3.9 5.3 6.0 6.7 9.3
NSR 3.2 4.1 6.5 8.6 10.4 10.9
AADNet 5.8 5.7 6.8 7.5 6.1 5.3
@@ 5.1 Classification performance
In line with most AAD studies, we started by training SS models and evaluating them using a multi-fold cross validation. However, we found unexpectedly low performances of the CCA, NSR, and AADNet models. While the exact reason for this is unclear, we conjecture that it is due to the higher number of parameters in these models. In consequence, the training requires a larger amount of data and the models are more susceptible to overfitting. This is particularly the case for the SS models and less of a problem for the SI models. To address this challenge, the SS models were trained by finetuning the pretrained SI model, achieving promising results. As shown in Figure 5, the proposed model outperformed the baseline methods on both datasets except for the DTU dataset with window lengths greater than 5 s where it achieved a comparable performance with the LSR method. This performance advantage could be attributed to two main factors: the nonlinearity and parallelized spatiotemporal convolutions, which allow the model to more accurately learn the audio representation in the human auditory system and capture the relationship between the audio stimulus and the brain signals.
@@
We also trained and evaluated SI models using the cross-trial LOSO validation scheme. Although all models yielded accuracies significantly above the chance performance, there was a consistent drop in performance compared to the corresponding SS models. This difference is somewhat expected, as the SS models, unlike the SI models, are fine-tuned to capture the unique characteristics of individual subjects. The results in Figure 6 showed a superior performance of the proposed AADNet compared to the other methods with a gap of 4.2 percentage points on the EventAAD dataset and 4.6 percentage points on the DTU dataset at the longest window length to the best baseline method. This demonstrated a better generalization for new subjects. We attribute this improvement to the multiple parallelized convolutions in the Inception structure which provides extensive coverage across subjects. Even though the improvement may seem modest, it could have a valuable contribution to the AAD field due to the training-free advantage of the SI model for the new subject. This advantage makes the SI model more feasible to be integrated into real-life applications. Furthermore, the limitation of SI models at 1 s window necessitates future work to investigate a more advanced model to generalize better at short window lengths.
@@
In regard to MESD, we showed that the proposed AADNet consistently and significantly obtained the lowest MESD values for both SS and SI models on both datasets. Since MESD represents the minimal expected time for an AAD-based gain control system to switch operation modes, these improvements demonstrated that the proposed AADNet holds a promise for integrating the AAD algorithm into a real-time gain control system of hearing-assistive devices.
@@ 5.2 Limitations
In this study, only four approaches were considered: a linear (LSR) and a non-linear (NSR) backward model, a forward-backward combined model (CCA), and the proposed direct classification model. We did not test the forward approach as it is an underperforming method [20, 21, 22]. Moreover, there have been several studies using neural networks and different features to solve different attention-decoding tasks, including speaker identification (SpkI) [9, 41, 13, 14, 15] and locus of attention (LoA) [42, 43, 44, 45]. It is challenging to make a direct comparison across these studies due to variations in datasets, used features, tasks, and analysis window lengths used to report results. Here, only the methods in the SpkI task that exploit the envelope-following response are included. For the sake of completeness and transparency, it must be mentioned that we also implemented the methods proposed by De Taillez et al.[13] and Ciccarelli et al.[14]. However, despite our best efforts in implementation and validation, these methods performed significantly worse than others, so their results are not included here. A similar observation was reported in [22].
@@ 5.3 Deep-learning methodology for direct AAD
The architecture of the network was inspired by the Inception block, which comprises multiple convolutional branches. The convolution in each branch acts as a spatiotemporal filter to capture how the speech is encoded in the neural response. The kernel size plays a role in limiting this relationship with a specific time delay. This structure allows the model to combine information in a way similar to the CCA method. However, in the CCA method, the optimal filter length may not work well for a wide range of subjects and datasets. With the parallel structures, in principle, we can add as many branches with different kernel sizes to cover a given available dataset as long as the dataset is sufficiently large for training, and therefore have better generalization.
@@
Another important design of the network in this study is the multi-audio input. This allows the network, at any time point during the training phase, to have access to all audio streams and have a higher degree of freedom to pick relevant features to maximize the difference between audio streams. This is different from the previous AAD studies where the models were forced to find the relationship between the brain response and the attended/unattended stimulus. Additionally, this structure allows us to straightforwardly extend the model to deal with more than two competing speakers and easily augment the data by shuffling the order of audio stimuli. This data augmentation, in turn, could eliminate the potential biases in directional attention, which are inevitable in some datasets, as pointed out in [34].

### v2

@@ 3.2.2 Audio
The audio envelopes were extracted using the compressed subband envelopes, resembling the processing of speech signals in the human auditory system [13]. Specifically, we first applied a gammatone filter bank with an equivalent rectangular bandwidth (ERB) equal to 1.5 Hz, and center frequencies ranging from 150 Hz to 4 kHz. The compressed envelope of each subband was computed with a power law exponent of 0.6. The final envelope was then obtained by summing all the subband envelopes. For the NSR model, the envelopes were extracted by taking the absolute value of the Hilbert transform of each speech stream. The envelopes were then low-pass filtered with a cut-off frequency of 50 Hz using an FIR filter, Hamming window, 12.5 Hz transition bandwidth, and resampled to 125 Hz. Both envelope extraction methods were implemented in Python using the SciPy package.
@@ 3.3 Evaluation procedure
As pointed out in Rotaru et al.[38], the DL approach is susceptible to capturing subtle biases, such as within-trial fingerprints of neural activities, even across different subjects. This may lead to artificially high decoding accuracies. To mitigate this potential bias, it is essential to perform appropriate data splitting into training, validation, and test sets. In this study, we carefully employ a trial-based cross validation to evaluate the SS and SI models for all investigated methods. In the remainder of this paper, we refer to the length of data used by the models to make a decision as the analysis window length.
@@ 3.3.2 SI models
For the SI model, data from one subject were held out for the test set while the data from the remaining subjects were used for training and validation. However, the EventAAD dataset was collected using the same pairs and order of stimuli for all subjects which could lead to potential leakage of stimuli information from the training set to the test set, allowing the model to indirectly learn to perform the AAD task. This is not the case for the DTU and KUL datasets where the order and pair of stimuli were randomly mixed. To ensure that the attended stimuli in the testing data were not presented as the attended stimuli in the training and validation data, we performed an eight-fold leave-one-subject-out (LOSO) cross-validation scheme. The data from the test subject were divided into eight folds. Only one fold was held out for the test set while the other seven folds were not used in the current iteration. For all the remaining training subjects, the trials whose attended stimuli being in the test set were left out of the training set, and the remaining trials were pooled together across subjects and split into a training set and a validation set in a 4:1 ratio (based on a trial basis). Models were trained using the training and validation set and evaluated on the test set to obtain a fold performance. In this manner, all the attended stimuli used for the test subject were not used as the attended stimuli during training, even for the other subjects. See Figure 4 for a visual illustration of the procedure. This procedure was repeated across eight folds. The accuracy of each SI model was calculated as the average accuracy across the eight folds.
@@ 3.4.2 Minimal expected switch duration
The accuracy metric described above is a window length-dependent quantification. It is expected that the longer the window length, i.e., the more information available, the higher the accuracy. To obtain a more effective performance measurement, we also calculated the minimal expected switch duration (MESD), an interpretable performance metric for AAD algorithms in the context of neuro-steered gain control [39]. The MESD addresses the trade-off between AAD accuracy and decision time by modeling an adaptive gain control system in a hearing device as a Markov chain, and based on that calculating the minimal expected time required to switch the operation mode after an attention switch of the user. A lower MESD corresponds to better AAD performance and vice versa. In this study, the MESD was calculated using the Python MESD toolbox [40].
@@ 3.5 Hyperparameter choice and model training
The hyper-parameters of the linear methods were selected based on the recommended values used in the original studies as discussed in Section 2. To determine the optimal $J$ value in the CCA method, we first performed inner cross-validation on the training data for the LDA model to find the optimal $J_{f}$ for each fold and each analysis window from the outer cross-validation of the CCA model. The final $J$ values for SS and SI models were obtained by taking the minimum value across analysis windows of the average and grand average of the $J_{f}$, respectively. These final $J$ values were then used to validate the corresponding CCA models again on the test data to obtain the final performances. For the NSR model, we performed full hyperparameter optimizations for each dataset based on the guidelines from the original study [17].
The proposed AADNet was trained with the Cross-Entropy loss function using the AdamW optimizer [41] which is advantageous in decoupling L2 regularization, i.e., weight decay, from the learning rate so that we have less number of hyper-parameters to search. The hyperparameters were chosen via a random search within configurations on the EventAAD dataset as follows: batch size = (32, 64, 128), weight decay = ($10^{-4}$, $10^{-3}$, $10^{-2}$, $10^{-1}$), dropout = (0.5, 0.4, 0.3, 0.2, 0.1) and number of output channel of FC1 layer, i.e., hidden units, $h$ = (32, 16, 0). In the case of $h$ = 0, the ELU, DO, BN, and FC2 after the FC1 layer were omitted. The learning rate was fixed at $5\times 10^{-5}$. The optimal hyperparameters were then fixed for the other two datasets. During the training process, the model was saved if the validation loss decreased and the training process was stopped if the validation loss did not decrease in at least 1 out of 5 consecutive epochs. Due to limited data sets for SS models that potentially lead to overfitting, for the NSR and the proposed AADNet models, we started training the SI models first and used the saved model to fine-tune the SS models for the subject that was left out with smaller tuned learning rates.
@@ 4.1 SS models
[…]
Figure 5: Comparison of the SS models on the three datasets. (a), (b) The accuracies and MESD of the four methods on the EventAAD dataset. (c), (d) The accuracies and MESD of the four methods on the DTU dataset. (e), (f) The accuracies and MESD of the four methods on the KUL dataset. The number of data points with an MESD of $>100\,$s is indicated as (+x) and was included in the computation of the medians. Statistical significance is denoted by asterisks (None: $p\geq 0.05$; *: $0.05>p\geq 0.01$; **: $0.01>p\geq 0.001$; ***: $0.001>p$.
Additionally, we calculated the per-subject MESD values for each model on the three datasets and presented them in Figure 5, Figure 5, and Figure 5, respectively. Significant differences in the median values between the proposed model and the others were also tested using paired permutation tests. We found that the proposed AADNet achieved a significantly lower MESD value of 21.9 s on the EventAAD dataset. It also achieved 12.3 s and 19.4 s on the DTU and KUL datasets, respectively. In comparison, the NSR model, consistent with the lowest performance in accuracy, achieved MEDS values of 63.1 s, 18.1 s, and 29.7 s for the EventAAD, DTU, and KUL datasets, respectively.
@@ 4.2 SI models
To evaluate how well AADNet generalizes to new subjects, we performed the SI validation scheme described in Section 3.3.2. The results of SI models are shown in Figure 6. Similar to the SS results, Figure 6, Figure 6, and Figure 6 present the test accuracies across subjects while Figure 6, Figure 6, and Figure 6 present the MESDs. Generally, the SI models obtained a lower performance compared to the corresponding SS models (see Table 3). The results demonstrate that AADNet significantly outperforms the other methods for almost all window lengths on all three datasets, except the CCA model at 1 s and the LSR model at 1 s (only on the KUL dataset). At this analysis window, AADNet and CCA models achieved similar performance, with accuracies of 0.563 and 0.561 for EventAAD, 0.575 and 0.577 for DTU, and 0.56 and 0.563 for KUL. The mean accuracies of AADNet reach 0.781 at 20 s, 0.894 at 40 s, and 0.826 at 40 s, for EventAAD, DTU, and KUL datasets, respectively. Regarding MESD, AADNet achieves significantly lower values of 27.8 s, 17.3 s, and 29.7 s on the EventAAD, DTU, and KUL datasets, respectively. The NSR method again performed the worst in both metrics.
Figure 6: Comparison of the SI models on the three datasets. (a), (b) The accuracies and MESD of the four methods on the EventAAD dataset. (c), (d) The accuracies and MESD of the four methods on the DTU dataset. (e), (f) The accuracies and MESD of the four methods on the KUL dataset.
Table 3: Performance drop (in percentage points of accuracy) of SI models compared to corresponding SS models.
@@
EventAAD LSR 1.6 1.8 2.5 3.5 3.7
CCA -0.3 0.5 -0.2 -0.4 -0.7
NSR 0.5 1.4 2.5 3.4 2.9
AADNet 0.9 1.5 2.3 3.0 4.1
DTU LSR 3.7 5.9 8.1 9.8 9.7 9.8
CCA 2.5 4.2 5.7 5.6 6.0 0.0
NSR 1.7 3.3 4.4 5.5 6.0 6.1
AADNet 2.2 3.6 5.1 6.1 6.4 5.5
KUL LSR 3.0 4.3 6.0 7.3 8.4 9.3
CCA 4.0 5.3 6.9 7.6 7.2 3.6
NSR 2.0 3.2 5.3 7.0 8.3 7.5
AADNet 1.9 2.9 4.2 5.7 6.1 6.8
@@ 5.1 Classification performance
In line with most AAD studies, we started by training SS models and evaluating them using a multi-fold cross-validation. However, we found unexpectedly low performances of the NSR and AADNet models. While the exact reason for this is unclear, we conjecture that it is due to the higher number of parameters in these models. In consequence, the training requires a larger amount of data and the models are more susceptible to overfitting. This is particularly the case for the SS models and less of a problem for the SI models. To address this challenge, the SS models were trained by finetuning the pretrained SI model, achieving promising results. As shown in Figure 5, the proposed model outperformed the baseline methods on the EventAAD dataset. Although it did not show superiority over the LSR method on the DTU dataset and the CCA method on the KUL dataset, AADNet achieved the best accuracy at long analysis windows($>$ 20 s) and an overall competitive performance on these two datasets, showcasing its strong adaptability across varying data distributions. This performance advantage could be attributed to two main factors: the nonlinearity and parallelized spatiotemporal convolutions, which allow the model to more effectively learn the audio representation in the human auditory system and capture the relationship between the audio stimulus and the brain signals.
@@
We also trained and evaluated SI models using the cross-trial LOSO validation scheme. Although all models yielded accuracies significantly above the chance performance, there was a consistent drop in performance compared to the corresponding SS models (see Table 3). This difference is somewhat expected, as the SS models, unlike the SI models, are fine-tuned to capture the unique characteristics of individual subjects. The results in Figure 6 showed a superior performance of the proposed AADNet compared to the other methods with gaps of 5.4, 5.9, and 4.0 of percentage points at the longest window length to the best baseline method on the EventAAD, DTU, and KUL datasets, respectively. This demonstrated a better generalization for new subjects. We conjecture that the multiple parallelized convolutions in the Inception structure, which provide extensive coverage of the spatiotemporal dependency between the speech envelope and the brain response across subjects, play a key role in this enhanced performance. Even though the improvement may seem modest, it could have a valuable contribution to the AAD field due to the training-free advantage of the SI model for the new subjects and the use of fixed hyperparameter sets across datasets. This advantage makes the SI model more feasible to be integrated into real-life applications.
@@
In regard to MESD, we showed that the proposed AADNet consistently and significantly achieved the lowest MESD values among SI models across all three datasets. Since MESD represents the minimal expected time for an AAD-based gain control system to switch operation modes, these improvements demonstrated that the proposed AADNet holds promise for integrating the AAD algorithm into a real-time gain control system for hearing-assistive devices. However, due to its moderate performance at short window lengths, the proposed SS model did not outperform the linear methods in MESD metrics. This motivates future work to develop a more advanced model to improve the performance at short window lengths.
5.2 Channel distribution of the AAD performance
To examine the importance of each channel to the overall performance of the AADNet, a leave-one-channel-out analysis was performed on the SI models for the three datasets. For each dataset, we trained and validated $N$ models in which one channel was left out at a time and the EEG data was referenced to the average of the remaining $N-1$ channels, where $N$ is the number of EEG electrodes used in that dataset. The performance drop of each model, compared to the original SI model, was then calculated and used to obtain the heatmaps (see the supplementary materials). The heatmaps consistently show a peak at the area of electrode C5 across the three datasets, indicating an active brain region involved in the speech attention task. This result is reasonable, as this area is close to the region of the primary auditory cortex and the superior temporal gyrus, which are primarily involved in speech comprehension and the envelope tracking process [45, 46]. Furthermore, the left lateralization of the heatmaps aligns with previous findings that language processing is strongly lateralized to the left hemisphere [47].
@@ 5.3 Limitations
In this study, only four approaches were considered: a linear (LSR) and a non-linear (NSR) backward model, a forward-backward combined model (CCA), and the proposed direct classification model. We did not test the forward approach as it is an underperforming method [22, 23, 24]. Moreover, there have been several studies using neural networks and different features to solve different attention-decoding tasks, including SpkI [11, 35, 15, 16, 17] and LoA [48, 49, 50, 51]. It is challenging to make a direct comparison across these studies due to variations in datasets, used features, tasks, and analysis window lengths used to report results. Here, only the methods in the SpkI task that exploit the envelope-following response are included. For the sake of completeness and transparency, it must be mentioned that we also implemented the methods proposed by De Taillez et al.[15], Ciccarelli et al.[16], and Kuruvila et al.[21]. However, despite our best efforts in implementation and validation, these methods performed significantly worse than others, so their results are not included here. A similar observation was reported in [24].
@@ 5.4 Deep-learning methodology for direct AAD
The architecture of the network was inspired by the Inception block, which comprises multiple convolutional branches. The convolution in each branch acts as a spatiotemporal filter to capture how the speech is encoded in the neural response. The kernel size plays a role in limiting this relationship with a specific time delay. This structure allows the model to combine information in a way similar to the CCA method. However, in the CCA method, the optimal filter length may not work well for a wide range of subjects and datasets. With the parallel architecture, the model can be extended by adding additional branches with different kernel sizes. This may enhance generalization performance, provided that a sufficient amount and diversity of training data are available.
@@
Another important design of the network in this study is the multi-audio input. This allows the network, at any time point during the training phase, to have access to all audio streams and have a higher degree of freedom to pick relevant features to maximize the difference between audio streams. This is different from the previous AAD studies where the models were forced to find the relationship between the brain response and the attended/unattended stimulus. Additionally, this structure allows us to easily augment the data by shuffling the order of audio stimuli. This data augmentation, in turn, could eliminate the potential biases in directional attention, which are inevitable in some datasets, as pointed out in [38].